# Uncovering Grounding IDs: How External Cues Shape Multi-Modal Binding

### v1

@@ Abstract
Large vision–language models (LVLMs) show strong performance across multimodal benchmarks but remain limited in structured reasoning and precise grounding. Recent work has demonstrated that adding simple visual structures, such as partitions and annotations, improves accuracy, yet the internal mechanisms underlying these gains remain unclear. We investigate this phenomenon and propose the concept of Grounding IDs, latent identifiers induced by external cues that bind objects to their designated partitions across modalities. Through representation analysis, we find that these identifiers emerge as robust within-partition alignment in embedding space and reduce the modality gap between image and text. Causal interventions further confirm that these identifiers mediate binding between objects and symbolic cues. We show that Grounding IDs strengthen attention between related components, which in turn improves cross-modal grounding and reduces hallucinations. Taken together, our results identify Grounding IDs as a key symbolic mechanism explaining how external cues enhance multimodal binding, offering both interpretability and practical improvements in robustness.
@@ 1 Introduction
Large vision–language models (LVLMs) such as LLaVA (Liu et al., 2023b), GPT-4V (Achiam et al., 2023), and Qwen-VL (Bai et al., 2023) have achieved strong results in multimodal tasks, including image captioning, visual question answering, and embodied tasks. Despite this progress, they remain limited in structured reasoning (Campbell et al., 2024), precise grounding, and hallucination (Liu et al., 2024). Very recent studies suggest that external visual structure can systematically improve LVLM reasoning. For instance, Rudman et al. (2025) reveals that many LVLMs struggle with basic polygon identification, while adding shape annotations substantially boosts accuracy. Notably, VISER (Izadi et al., 2025) introduces more general visual structure (e.g., grids or horizontal lines) and demonstrates consistent gains across diverse reasoning tasks. Despite compelling advancements, the interpretability and mechanism-level understanding of this phenomenon remain underexplored. Prior work has largely focused on what visual structure can enhance reasoning, but not how it helps.
@@ 2 Background and Hypotheses
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that links entities with their attributes (Feng & Steinhardt, 2023). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they employ similar vectors linking visual objects with textual references (Saravanan et al., 2025), with properties such as factorizability and position independence. Inspired by the notion of such abstract identifiers, we investigate the underlying circuit by which external visual structures equipped with textual cues enhance LVLMs’ reasoning.
@@ 3 Empirical Evidence for Existence of Grounding IDs
To investigate why external cues improve performance, we first analyze internal attention patterns of LVLMs. We construct a controlled synthetic dataset where each image contains 15 unique objects, sampled from 35 shape–color combinations (7 shapes $\times$ 5 colors). The model is tasked with generating scene descriptions. The model is tasked with generating scene descriptions. As external cues, we add symbols (&, #, \$, @) to both the image and the prompt, and separate the image into four partitions using three horizontal lines (Fig. 2). Ablations on alternative cue designs are provided in the appendix. To analyze embeddings and attention, each token must be paired with its counterpart from the other modality and assigned to the corresponding partition. Visual tokens are grouped according to their position within the four rows. Text tokens are mapped to visual objects using regular expressions, aided by the fact that objects are unique within each image. This yields a partition pseudo-label for every visual and textual token, indicating the partition to which it belongs.
Figure 2: Illustration of attention patterns under baseline and structured inputs. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices. Values are averaged over 100 samples and layers 22–27 of Qwen2.5-VL (Bai et al., 2025).
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, increasing alignment and grounding. We now seek causal validation. For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols (&, #, \$, @) to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, ”What is in row ?”, and the model must respond ”red circle.” To formalize notation, let us define $\textbf{o}^{s_{j}}_{s_{i}}$ as an object o that is located in the partition associated with symbol $s_{i}$ and bound to symbol $s_{j}$. For example, $\textbf{o}^{\mathdollar}_{\&}$ denotes an object positioned in the & partition but bound to \$. For convenience, we use two shorthand cases: (1) $\textbf{o}^{\sim s}_{s}$ for an object located in the partition of $s$ but bound to a different symbol, and (2) $\textbf{o}^{s}_{\sim s}$ for an object bound to $s$ but located in a different partition.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to prior studies (Vig et al., 2020; Feng & Steinhardt, 2023). Two images $c$ and $c^{\prime}$ are randomly sampled, and their activations are extracted across all layers. We then select two random symbols (e.g., \$ and #) and swap their corresponding object patch activations between the two contexts. For instance, the object in row \$ of $c^{\prime}$ is inserted into row # of $c$, while the object in row # of $c^{\prime}$ is patched into row \$. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are the same between $c$ and $c^{\prime}$. We intentionally choose the swap strategy to avoid ambiguity arising from the co-existence of $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{s}_{s}$.
We find that predictions consistently follow the binding symbols of the swapped objects rather than the symbols present locally in the host context. In other words, for a given partition $s$, the model outputs $\textbf{o}^{s}_{\sim s}$ instead of $\textbf{o}^{\sim s}_{s}$. For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @ even though it is located beside &. This indicates that the model disregards local cues and follows the transferred symbol–object binding (Fig. 4(a)).
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ across layers. (b) Layer- and head-wise signal-to-noise of attention differences between $\textbf{o}^{s}{\sim s}$ and $\textbf{o}^{\sim s}{s}$.
Next, we investigate in which layers the model begins to predict the intervened grounded object ($\textbf{o}^{s}_{\sim s}$) rather than the adjacent object ($\textbf{o}^{\sim s}_{s}$). We apply the logit lens technique (nostalgebraist, 2020) to the single-word response token (using a monochrome dataset for simplicity). By computing the logit difference between the $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ and plotting averaged values across layers (Fig. 5(a)), we find that the model shifts toward predicting the intervened grounded object in the later layers (20–27). We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\textbf{o}^{s}_{\sim s}$ versus $\textbf{o}^{\sim s}_{s}$ across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers grounded objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in layer 16, which attend more to grounded objects.
@@ 5 Implications of Grounding IDs: Hallucination Reduction
Figure 7: Cross-attention behavior on the synthetic dataset with 10, 20, and 30 objects, averaged over 100 samples. Attention is computed using a max operation with a window size of 5 and stride of 1 across generated tokens.
Table 1: Evaluation on synthetic datasets containing 100 samples each with 10, 20, or 30 unique objects.
@@
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (Favero et al., 2024; Huang et al., 2024). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, measuring accuracy, precision, recall, and F1 for object descriptions, as reported in Table 1. Structured inputs consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens with increasing object counts. Notably, external cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves the faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (Lin et al., 2014), with performance assessed by the hallucination-specific metric CHAIR (Rohrbach et al., 2018). We conduct experiments on two recent LVLMs, Qwen2.5-VL (Bai et al., 2025), and LLaVA-1.5 (Liu et al., 2023a). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding. Results (Fig. 8 and Table 2) show substantial reductions in both CHAIR-s and CHAIR-i compared to baseline. Importantly, unlike most existing techniques, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (Hurst et al., 2024) and Gemini-2.5-Pro (Comanici et al., 2025), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and Sparc (Jung et al., 2025), while requiring no additional inference modules and maintaining near-zero computational overhead. We also include some failure cases and report results on another commonly used benchmark, POPE (Li et al., 2023), in the appendix.
@@
Table 2: Evaluation on MS-COCO using CHAIR metrics across open- and closed-source models. Results are reported for sentence-level (CHAIRs) and instance-level (CHAIRi) hallucination rates (lower is better). Inference time is shown where available.
@@ A.4 Evaluation on Synthetic Variants
  • Grid 4×4: A different baseline layout where objects are arranged in a $4\times 4$ grid. Both numbered and unlined versions were tested. Performance remains weaker than the structured 1D-row settings, highlighting that the grid layout introduces ambiguity without strong row-level cues.
@@ A.5 Implementation Details and Qualitative Example
Appendix B Evaluation Metrics for Object Hallucination
@@
B.0.1 CHAIR: Caption Hallucination Assessment with Image Relevance
@@
B.0.2 POPE: Probing Object-Presence Hallucination
@@
B.1 Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix B, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@ B.1 Evaluation on POPE Benchmark
Appendix C Activation swapping (interchange intervention)
@@
C.1 Disjoint-symbol swap (no-visual-cue prompt)

### v2

@@ Abstract
Large vision–language models (LVLMs) show strong performance across multimodal benchmarks but remain limited in structured reasoning and precise grounding. Recent work has demonstrated that adding simple visual structures, such as partitions and annotations, improves accuracy, yet the internal mechanisms underlying these gains remain unclear. We investigate this phenomenon and propose the concept of Grounding IDs, latent identifiers induced by external cues that bind objects to their designated partitions across modalities. Through representation analysis, we find that these identifiers emerge as robust within-partition alignment in embedding space and reduce the modality gap between image and text. Causal interventions further confirm that these identifiers mediate binding between objects and symbolic cues. We show that Grounding IDs strengthen attention between related components, which in turn improves cross-modal grounding and reduces hallucinations. Taken together, our results identify Grounding IDs as a key symbolic mechanism that explains how external cues enhance multimodal binding and offers both interpretability and practical improvements in robustness.
@@ 1 Introduction
Large vision–language models (LVLMs) such as LLaVA (Liu et al., 2023b), GPT-4V (Achiam et al., 2023), and Qwen-VL (Bai et al., 2023) have achieved strong results in multimodal tasks, including image captioning, visual question answering, and embodied tasks. Despite this progress, they remain limited in structured reasoning (Campbell et al., 2024), precise grounding, and hallucination (Liu et al., 2024). Very recent studies suggest that external visual structure can systematically improve LVLM reasoning. For instance, Rudman et al. (2025) reveal that many LVLMs struggle with basic polygon identification, while adding shape annotations substantially boosts accuracy. Notably, VISER (Izadi et al., 2025) introduces more general visual structure (e.g., grids or horizontal lines) and demonstrates consistent gains across diverse reasoning tasks. Despite compelling advancements, the interpretability and mechanism-level understanding of this phenomenon remain underexplored. Prior work has largely focused on what visual structure can enhance reasoning, but not how it helps.
@@ 2 Background and Hypotheses
Recent advances in mechanistic interpretability show that LLMs solve entity–attribute binding using Binding IDs, latent vectors that link entities with their attributes (Feng & Steinhardt, 2023). Causal mediation analyses support this finding and further show that these identifiers behave additively in representation space. Follow-up work extends this idea to vision–language models, showing that they employ similar vectors linking visual objects with textual references (Saravanan et al., 2025), with properties such as factorizability and position independence. Inspired by the notion of such abstract identifiers, we investigate the underlying circuit by which external visual structures equipped with textual cues enhance LVLMs’ reasoning.
@@ 3 Empirical Evidence for Existence of Grounding IDs
To investigate why external cues improve performance, we first analyze internal attention patterns of LVLMs. We construct a controlled synthetic dataset where each image contains 15 unique objects, sampled from 35 shape–color combinations (7 shapes $\times$ 5 colors). The model, Qwen2.5-VL (Bai et al., 2025), is tasked with generating scene descriptions. As external cues, we add symbols (&, #, \$, @) to both the image and the prompt, and separate the image into four partitions using three horizontal lines (Fig. 2). Ablations on alternative cue designs are provided in the appendix. To analyze embeddings and attention, each token must be paired with its counterpart from the other modality and assigned to the corresponding partition. Visual tokens are grouped according to their position within the four rows. Text tokens are mapped to visual objects using regular expressions, aided by the fact that objects are unique within each image. This yields a partition pseudo-label for every visual and textual token, indicating the partition to which it belongs.
Figure 2: Illustration of attention patterns under baseline and structured inputs. (a) One dataset sample (top: baseline, bottom: structured). (b) Within-modality visual attention matrices. (c) Cross-modality attention matrices. Values are averaged over 100 samples and layers 22–27 of Qwen2.5-VL.
@@ 4 Causal Evidence of Grounding IDs
The alignment and attention analyses in Section 3 provide correlational evidence that external cues induce partition-based alignment. Our hypothesis is that Grounding IDs act as abstract vectors that are induced to related tokens across modalities, increasing alignment and grounding. We now seek causal validation. For causal mediation analysis, we use a simplified synthetic dataset in which each image contains four rows with one object per row. Rows are labeled with non-ordinal symbols (&, #, \$, @) to prevent sequential order cues. The task is discriminative visual question answering: the prompt asks, for example, ”What is in row @?”, and the model must respond ”red circle.” To formalize notation, let us define $\textbf{o}^{s_{j}}_{s_{i}}$ as an object o that is located in the partition associated with symbol $s_{i}$ and bound to symbol $s_{j}$. For example, $\textbf{o}^{\mathdollar}_{\&}$ denotes an object positioned in the & partition but bound to \$. For convenience, we use two shorthand cases: (1) $\textbf{o}^{\sim s}_{s}$ for an object located in the partition of $s$ but bound to a different symbol, and (2) $\textbf{o}^{s}_{\sim s}$ for an object bound to $s$ but located in a different partition.
@@ 4.1 Activation Swapping Experiment
We follow a causal mediation framework similar to prior studies (Vig et al., 2020; Feng & Steinhardt, 2023). Two images $c$ and $c^{\prime}$ are randomly sampled, and their activations are extracted across all layers. We then select two random symbols (e.g., \$ and #) and swap their corresponding object patch activations between the two contexts. For instance, the object in row \$ of $c^{\prime}$ is inserted into row # of $c$, while the object in row # of $c^{\prime}$ is patched into row \$. We pass the patched input $c^{*}$ to the model and record its predictions for the selected symbols. Fig. 4(a) shows a simple case where the swapped objects are the same between $c$ and $c^{\prime}$. We intentionally choose the swap strategy to avoid ambiguity arising from the coexistence of $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{s}_{s}$.
We find that predictions consistently follow the binding symbols of the swapped objects rather than the symbols present locally in the host context. In other words, for a given partition $s$, the model outputs $\textbf{o}^{s}_{\sim s}$ instead of $\textbf{o}^{\sim s}_{s}$. For example, if the object in row @ of $c^{\prime}$ is a green square and it is inserted into row & of $c$, the model reports a green square for row @ even though that object is located beside &. This indicates that the model disregards local cues and follows the transferred symbol–object binding (Fig. 4(a)).
@@
Figure 5: Causal mediation analysis across layers. (a) Average logit differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ across layers. (b) Layer- and head-wise signal-to-noise of attention differences between $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ patches.
Next, we investigate in which layers the model begins to predict the intervened grounded object ($\textbf{o}^{s}_{\sim s}$) rather than the adjacent object ($\textbf{o}^{\sim s}_{s}$). We apply the logit lens technique (nostalgebraist, 2020) to the single-word response token (using a monochrome dataset for simplicity). By computing the logit difference between the $\textbf{o}^{s}_{\sim s}$ and $\textbf{o}^{\sim s}_{s}$ and plotting averaged values across layers (Fig. 5(a)), we find that the model shifts toward predicting the intervened grounded object in the later layers (20–27). We also investigate which attention heads are most responsible for propagating Grounding IDs as opposed to relying on local visual proximity. For each head and layer, we compute the difference in attention from the response token to visual tokens corresponding to $\textbf{o}^{s}_{\sim s}$ versus $\textbf{o}^{\sim s}_{s}$ across samples. To quantify consistency, we divide the mean difference by its standard deviation, yielding a signal-to-noise ratio that reflects how reliably a head prefers grounded objects. To emphasize attention to meaningful regions, this ratio is multiplied by the head’s average attention weight. The resulting scores are visualized in Fig. 5(b), showing that Grounding IDs cause significant attention shifts in certain heads, particularly in layer 16, which attends more to grounded objects.
@@ 5 Implications of Grounding IDs: Hallucination Reduction
Figure 7: Cross-attention behavior on the synthetic dataset with 20 objects, averaged over 100 samples. Attention is computed using a max operation with a window size of 5 and stride of 1 across generated tokens.
Table 1: Evaluation of the scene description task on synthetic datasets, each containing 100 samples with 10, 15, or 20 unique objects per image.
@@
Prior work attributes hallucinations in LVLMs to the decay of visual attention and increasing reliance on language priors (Favero et al., 2024; Huang et al., 2024). Motivated by this, we evaluate hallucination on the controlled datasets from Section 3, measuring accuracy, precision, recall, and F1 for object descriptions, as reported in Table 1. Structured inputs consistently improve performance across all metrics, with particularly strong gains in precision, which is most relevant for avoiding mentions of nonexistent objects. For images with 10 objects, baseline recall is slightly higher, but as the number of objects increases, structured inputs outperform the baseline across all criteria. The performance gap due to external structure also widens with increasing object counts. Notably, multimodal cues are considerably more effective than unimodal cues. These results highlight that sustained cross-modal grounding directly improves faithfulness of generated text.
Finally, we evaluate on large-scale hallucination benchmarks using MS-COCO images (Lin et al., 2014), with performance assessed by the hallucination-specific metric CHAIR (Rohrbach et al., 2018). We conduct experiments on two recent LVLMs, Qwen2.5-VL (Bai et al., 2025), and LLaVA-1.5 (Liu et al., 2023a). We observe that for natural scenes, simple horizontal lines are insufficient, and grid-based partitions provide more effective structure for sequential scanning of the image. To ensure visibility against diverse backgrounds, we add thin white margins around the scaffolding. Results (Fig. 8 and Table 2) show substantial reductions in both CHAIRs and CHAIRi compared to baseline. Importantly, unlike most existing techniques, the proposed approach is also applicable to black-box, closed-source models such as GPT-4o (Hurst et al., 2024) and Gemini-2.5-Pro (Comanici et al., 2025), which are considered strong models for visual reasoning. Notably, our simple strategy outperforms or matches specialized hallucination-mitigation methods such as VCD (Leng et al., 2024), OPERA (Huang et al., 2024), and SPARC (Jung et al., 2025), while requiring no additional inference modules and maintaining near-zero computational overhead. We also report results on another commonly used benchmark, POPE (Li et al., 2023), in the appendix.
@@
Table 2: Evaluation on MS-COCO using CHAIR metrics across open- and closed-source models. Results are reported for sentence-level (CHAIRS) and instance-level (CHAIRI) hallucination rates (lower is better). The average inference time measured on a machine with an NVIDIA RTX 4090 GPU.
@@ A.4 Evaluation on Synthetic Variants
  • Grid 4×4: A baseline layout where objects are arranged in a $4\times 4$ grid. Two variants were tested: numbered with grid lines, and numbered without grid lines. Performance remained weaker than in the structured row version.
@@ A.5 Implementation Details and Qualitative Example
A.6 Evaluation Metrics for Object Hallucination
@@
A.6.1 CHAIR: Caption Hallucination Assessment with Image Relevance
@@
A.6.2 POPE: Probing Object-Presence Hallucination
@@
A.6.3 Evaluation on POPE Benchmark
In addition to CHAIR, we evaluate our structured scaffolding method using the POPE benchmark, which probes object hallucination through structured binary questions. As described in Appendix A.6, POPE defines three subsets—Random, Adversarial, and Popular—each targeting different hallucination biases such as language priors, scene co-occurrence, and object frequency.
@@ A.6.3 Evaluation on POPE Benchmark
A.7 Activation swapping (interchange intervention)
@@
A.7.1 Disjoint-symbol swap (no-visual-cue prompt)